Mary can purchase roses and daisies from a flower shop. She can buy four roses and five daisies for $7.90 or two roses and six daisies for $5.70. what is the individual cost for each rose and daisy.
Answer by mananth(16949)

rose cost $ x
Daisy cost $ y

4x+5y=7.9 .............1
2x+6y=5.7 .............2

multiply (1)by -6
Multiply (2) by 5
-24x-30y=-47.4
10x+30y=28.5
Add the two equations
-14x=-18.9
/-14
x=1.35
plug value of x in (1)
4x+5y=7.9
5.4+5y= 7.9
5y=7.9-5.4
5y=2.5
y=0.50